OpenStudy (tainted):

Solving Quadratics by Factoring Solve x^2 = 14x – 49 (Show work for medal)

OpenStudy (campbell_st):

rewrite it x^2 - 14x + 49 = 0 its a perfect square so it should be easy to factorise... 2 factors of 49 that that are both negative and add to -14

OpenStudy (tainted):

@campbell_st x2 = 14x – 49 = x2 - 14x + 49 =0 = (x-7)^2=0 = x=7?

OpenStudy (campbell_st):

yep... its that simple...

OpenStudy (tainted):

I thought it'd be harder....Ha